OpenStudy (anonymous):

Given the function f(x) = 5(x+4) − 6, solve for the inverse function when x = 19.

OpenStudy (zzr0ck3r):

y=5(x+4)-6 switch the x,y and solve for y x=5(y+4)-6 now solve for y

OpenStudy (zzr0ck3r):

or 5(x+4)-6=19 solve for x here

OpenStudy (zzr0ck3r):

there are two ways to go about this....

OpenStudy (anonymous):

i distribute the 5 in the parentheses right?

OpenStudy (zzr0ck3r):

5x+20-6=19 5x+14=19 5x=5 x=?

OpenStudy (anonymous):

1!! :)

OpenStudy (zzr0ck3r):

correct